### v2

# GraphIC: A Graph-Based In-Context Example Retrieval Model for Multi-Step Reasoning

Abstract
In-context learning (ICL) enhances large language models (LLMs) by incorporating demonstration examples, yet its effectiveness heavily depends on the quality of selected examples. Current methods typically use text embeddings to measure semantic similarity, which often introduces bias in multi-step reasoning tasks. This occurs because text embeddings contain irrelevant semantic information and lack deeper reasoning structures. To address this, we propose GraphIC, a graph-based retrieval model that leverages reasoning-aware representation and specialized similarity metric for in-context example retrieval. GraphIC first constructs thought graphs—directed, node-attributed graphs that explicitly model reasoning steps and their dependencies—for candidate examples and queries. This approach filters out superficial semantics while preserving essential reasoning processes. Next, GraphIC retrieves examples using a novel similarity metric tailored for these graphs, capturing sequential reasoning patterns and asymmetry between examples. Comprehensive evaluations across mathematical reasoning, code generation, and logical reasoning tasks demonstrate that GraphIC outperforms 10 baseline methods. Our results highlight the importance of reasoning-aware retrieval in ICL, offering a robust solution for enhancing LLM performance in multi-step reasoning scenarios.
GraphIC: A Graph-Based In-Context Example Retrieval Model for Multi-Step Reasoning
Jiale Fu1, Yaqing Wang2,*, Simeng Han3, Jiaming Fan1, Xu Yang1,* 1Southeast University, 2Beijing Institute of Mathematical Sciences and Applications, 3Yale University, *Corresponding author.
1 Introduction
In-context learning (ICL) Brown et al. (2020) allows large language models (LLMs) to adapt to new tasks by incorporating a few demonstration examples within the input prompt, without updating model parameters. However, studies reveal that ICL performance heavily depends on the quality of selected in-context examples (ICEs) Zhao et al. (2021); Liu et al. (2022), motivating extensive research on ICE selection strategies. Current approaches Liu et al. (2022); Rubin et al. (2022) typically use text embeddings to measure semantic similarity between queries and candidate examples, achieving success in semantic-centric tasks like text classification and translation Agrawal et al. (2023).
However, these text-based methods face significant limitations in multi-step mathematical and logical reasoning tasks. This is because text embedding encodes a substantial amount of shallow semantic information, which is irrelevant to the underlying reasoning processes. This extraneous information introduces bias in the selection of ICEs An et al. (2023). As shown in Figure 1 (left), when solving a speed calculation problem, text-based methods may retrieve examples about distance/time calculations due to shallow semantic similarity, inducing incorrect reasoning paths (e.g., calculating distance instead of speed). This observation motivates our key insight: Effective ICE selection for reasoning tasks requires representations that explicitly model cognitive processes rather than textual surface forms.
Drawing from cognitive science Friston (2008) and graph-based reasoning works Besta et al. (2024); Yao et al. (2023), we propose thought graphs—directed node-attributed graphs where nodes represent reasoning steps and edges denote step dependencies (i.e., a child step can only proceed after the parent step is completed). This representation filters irrelevant semantics while preserving essential reasoning patterns. An example of a thought graph is shown in Figure 2 (a).
Figure 1: ICL with different ICE retrieval mechanisms. The left panel shows examples retrieved via BERT embedding Devlin et al. (2019), while the right panel displays examples retrieved via GraphIC. Semantically related terms are highlighted in blue, and quantities or variables needing resolution are highlighted in green.
Furthermore, we propose a similarity metric for thought graphs tailored to multi-step reasoning tasks. Specifically, we introduce a parameter $W$ to model the reasoning pattern of a given process. Given two thought graphs, $G_{1}$ and $G_{2}$, we first estimate the reasoning pattern parameter $W_{1}$ for $G_{1}$, then assess its applicability to $G_{2}$. This applicability serves as the similarity measure $s(G_{1},G_{2})$. Our approach naturally captures the sequential structure of reasoning steps and accounts for the inherent asymmetry between examples, which we discuss in detail in Section 3.2.
Building upon these foundations, we introduce GraphIC, a Graph-based In-Context Example Retrieval Model. GraphIC achieves reasoning-aware example selection through three key phases: (1) constructing thought graphs for both the query and candidate examples, (2) calculating graph similarity using our proposed metric, and (3) retrieving the top-$k$ most relevant examples. As illustrated in Figure 1 (right), GraphIC effectively identifies examples that align with the reasoning process (such as calculations for the number of envelopes to stuff per hour or unit price), even if they lack semantic similarity to the query (e.g., speed computation). This enables the LLM to solve the problem correctly.
Through comprehensive evaluations across mathematical reasoning, code generation, and logical reasoning tasks, GraphIC demonstrates superior performance over 10 baseline methods including both training-free and training-based approaches. To sum up, our key contributions are: (1) A representaion.We introduce a novel graph-based representation, called thought graph, to model multi-step reasoning processes. This representation effectively filters out irrelevant shallow semantic information while preserving the essential reasoning steps. (2) A similarity metric.We introduce a similarity metric for thought graphs tailored to multi-step reasoning tasks that capture the sequential nature of the steps and the asymmetry between examples. (3) Empirical validation.Our experimental results indicate that GraphIC, despite being a training-free model, outperforms both training-free and training-based models across various multi-step reasoning tasks.
2 Related Work
Existing ICE selection techniques can be classified as either training-free or training-based, depending on whether a retriever needs to be trained.
Training-free approaches are generally divided into two types: (i) those that use heuristic criteria such as similarity Liu et al. (2022); Hu et al. (2022), diversity Cho et al. (2023); Zhang et al. (2022b); Levy et al. (2023); Hongjin et al. (2022); Zhang et al. (2023), complexity Fu et al. (2022), or combinations of these Agrawal et al. (2023); Tonglet et al. (2023); Gupta et al. (2023) to select in-context examples (ICEs); (ii) those that leverage feedback from LLMs, such as probability distributions Wu et al. (2023); Nguyen and Wong (2023); Li and Qiu (2023); Yang et al. (2023), perplexity Gonen et al. (2023), or the model’s generated output An et al. (2023) to guide the selection process. Training-free approaches eliminate the computational and time costs of model training, but their simpler architecture often leads to lower performance than training-based methods.
Training-based methods are generally divided into two types. The first learns to select individual examples and then extends this to $k$-shot scenarios Rubin et al. (2022); Xiong et al. (2024); Gupta et al. (2024). The second models the selection of a group of examples as a whole Ye et al. (2023); Wang et al. (2023); Zhang et al. (2022a); Scarlatos and Lan (2023); Lu et al. (2022); Peng et al. (2023); Xu et al. (2024). Training-based approaches often deliver better performance, but they are computationally expensive and time-consuming.
Our proposed GraphIC involves using LLM-generated output to select ICEs, similar to Skill-kNN and DQ-LoRe. Although methods increase retrieval time due to invoking the LLM during the retrieval, their use of LLMs makes them suitable for more complex tasks.
3 The Proposed GraphIC
3.1 Thought Graphs
Figure 2: An example of a thought graph (a) and its corresponding FRR (b).
Introduction of Thought Graphs. As outlined earlier, thought graphs are directed, node-attributed graphs designed to model the reasoning process explicitly. In a thought graph, each node $i$ is attributed with an embedding vector $x_{i}\in\mathbb{R}^{d}$ of a text, representing the key information of the current reasoning step. Directed edges between vertices represent the dependencies between steps. For example, if step B depends on the completion of step A, there is an edge from step A to step B.
[…]
Due to the inherent differences between tasks, the text used to compute embedding varies slightly. Specifically, for mathematical reasoning, code generation, and logical reasoning tasks, we use mathematical operations, code snippets, and intermediate conclusions, respectively, as input texts. This is because, in mathematical reasoning tasks, a reasoning step can be represented by "which mathematical operations were performed on which variables"; in code generation tasks, it corresponds to "what code was executed"; and in logical reasoning tasks, it reflects "which existing conclusions were used to derive new conclusions". Examples of thought graphs for the four benchmark datasets are provided in Appendix A.1.
Construction of Thought Graphs. For candidate examples, which include ground-truth natural language reasoning processes (i.e., chain-of-thought answers), we employ the following method to generate thought graphs. First, we use an LLM to formalize the reasoning process for each example. This formalized representation is called the Formalized Reasoning Representation (FRR). FRR filters irrelevant shallow semantics and extracts the key information from the reasoning process. Moreover, FRR is designed to be easily converted to a graph by applying rule-based parsing. The prompts used for generating FRRs and the pseudocode for parsing FRRs are shown in Appendix A.2.
Figure 2 (b) illustrates the FRR of a thought graph. Each line in the FRR follows the format A = [B](C), where A represents the name of the vertex, denoting the vertex itself. B represents the label of the vertex, which will be used to compute its embedding. C denotes the parent vertex of the given vertex. For example, in the FRR shown in Figure 2 (b), by parsing the first line, we create node_1, set its label to "add", and add edges from "20" and "0.3" to node_1. The remaining lines are parsed similarly to the first one. Note that in Figure 2 (a), the vertices representing numbers are indicated by dashed lines, which signifies that these vertices do not represent specific operations and will be removed once the entire graph is constructed.
For queries, since they lack ground-truth reasoning processes, we first use an LLM to generate pseudo natural language reasoning processes. We then follow the same procedure as with candidate examples to create thought graphs. Note that it is unnecessary to ensure the correctness of the generated reasoning process, as it is only used to aid in the retrieval of ICEs and is not part of the final answer. In fact, when a generated answer is incorrect, it is often not entirely wrong, but partially correct, which still reflects the reasoning process and can be helpful for retrieving examples. This ensures that GraphIC is robust to inaccuracies in the generated reasoning process. Even when the reasoning process is incorrect, GraphIC still maintains high accuracy. A detailed discussion and related experiments are provided in Appendix A.3.
3.2 Similarity Measure for Thought Graphs
In this subsection, we first introduce a simplified case to illustrate our core idea, then extend it to propose our similarity measure.
Simplified Case. Consider a thought graph $G$ with vertices $v_{1},\dots,v_{n}$, where each vertex $v_{i}$ corresponds to a reasoning step with an attribute vector $x_{i}\in\mathbb{R}^{d}$ (e.g., an embedding of a mathematical operation). Let $X=[x_{1},\dots,x_{n}]\in\mathbb{R}^{n\times d}$ denote the matrix of concatenated attributes. The adjacency matrix $A$ of $G$ is defined such that $A_{ij}=1$ if there is a directed edge from $v_{i}$ to $v_{j}$, and $A_{ij}=0$ otherwise. For a vertex $v_{i}$, let $\text{pa}(v_{i})$ denote the set of its parent vertices (direct predecessors).
Generally, given two thought graphs $G_{1}$ and $G_{2}$, the similarity $s(G_{1},G_{2})$ is computed in two steps: (1). Extract Reasoning Pattern: Solve Equation (1) for $G_{1}$ to obtain $W_{1}$, which encapsulates $G_{1}$’s reasoning pattern. (2). Compute Applicability: Evaluate how well the reasoning pattern $W_{1}$ applies to thought graph $G_{2}$. This applicability, denoted as $\mathcal{A}(W_{1};G_{2})$, serves as the similarity measure $s(G_{1},G_{2})$. We now provide a detailed explanation of these two steps:
Firstly, we derive a matrix $W$ that captures directional relationships in $G$. This is achieved by solving the following optimization problem:
$W$ $=\mathop{\arg\max}\limits_{\begin{subarray}{c}W\in\mathbb{R}^{d\times d},\\ \|W\|_{F}=1\end{subarray}}\mathcal{A}(W;G),\quad\text{where}$ (1)
$\mathcal{A}(W;G)$ $=\sum_{i=1}^{n}\underbrace{(\sum_{j\in\text{pa}(v_{i})}x_{j})^{\top}}_{z_{i}}Wx_{i}.$
Here, the term $z_{i}=\sum_{j\in\text{pa}(v_{i})}x_{j}$ calculates the sum of the attributes of $v_{i}$’s parent nodes, aggregating precursor information for step $v_{i}$. $z_{i}^{\top}W$ represents a linear transformation of this information. $z_{i}^{\top}Wx_{i}$ measures the inner product similarity between transformed precursor information $Wz_{i}$ and the current step’s attribute $x_{i}$. By maximizing $\mathcal{A}(W;G)$ under the Frobenius norm constraint ($\|W\|_{F}=1$), $W$ learns to predict the attribute of the next step based on precursor information, effectively encoding the directional relationships in $G$.
For instance, consider a thought graph $G$ with two vertices: $v_{1}$ with attribute $x_{1}=\operatorname{Emb(\text{"add"})}$ and $v_{2}$ with $x_{2}=\operatorname{Emb(\text{"multiply"})}$, connected by an edge $v_{1}\to v_{2}$. In this setup, $W$ learns to map $x_{1}$ ("add") to $x_{2}$ ("multiply"), indicating that "multiply" typically follows "add" in $G$. When this learned mapping is applied to another graph $G^{\prime}$, a high value of $\mathcal{A}(W;G^{\prime})$ suggests that $G^{\prime}$ also contains frequent "add" $\to$ "multiply" sequences, reflecting similar problem-solving logic. Therefore, $\mathcal{A}(W;G^{\prime})$ quantifies how well the reasoning pattern of $G$ applies to $G^{\prime}$. A higher $\mathcal{A}(W;G^{\prime})$ indicates an alignment between the two graphs’ transitions, indicating that $G^{\prime}$ follows a similar problem-solving trajectory. This metric, therefore, provides an effective measure of the similarity in reasoning between $G$ and $G^{\prime}$, making it suitable for retrieving ICEs that share compatible problem-solving logic.
Proposed Similarity Metric. The similarity metric in GraphIC builds upon the simplified case while introducing two critical enhancements to better model reasoning dynamics and computational efficiency.
Firstly, in Equation (1), the information state $z_{i}$ for each node aggregates only its direct parent attributes. However, human reasoning often incorporates information from multiple prior steps and occasionally revisits initial premises. To capture this, we propose an iterative aggregation process:
$Z^{(h+1)}=[(1-\lambda)\tilde{A}+\lambda\tilde{B}+I]Z^{(h)},$ (2)
where $Z^{(h)}$ represents node information states at iteration $h$; $Z^{(0)}$ is initialized with node attributes $X$; $I$ is identity matrix; $\tilde{A}=D_{A}^{-\frac{1}{2}}AD_{A}^{-\frac{1}{2}}$, $D_{A}=\operatorname{diag}(\operatorname{deg}_{\operatorname{out}}(v_{1}),\dots,\operatorname{deg}_{\operatorname{out}}(v_{1}))$; $\tilde{B}=D_{B}^{-\frac{1}{2}}BD_{B}^{-\frac{1}{2}}$, and $B$ is traceback matrix enabling backtracking to root nodes, defined as $B_{ij}=1$ if $\text{deg}_{\text{in}}(v_{j})=0$ and $\text{deg}_{\text{in}}(v_{i})>0$, otherwise $B_{ij}=0$, with $\text{deg}_{\text{in}}(v_{j})$ representing the in-degree of node $v_{j}$. A visual example of matrix $B$ is presented in Appendix A.4. $\lambda$ is a balancing hyperparameter ($0\leq\lambda\leq 1$).
In each iteration, $z_{i}$ is updated by combining three information sources, corresponding to the matrices $A$, $B$, and $I$. Specifically, $AZ^{(h)}$ aggregates information from direct parent nodes, similar to the simplified case; $BZ^{(h)}$ facilitates backtracking in the reasoning process; and $IZ^{(h)}$ maintains current information.
Secondly, we observe that directly solving the optimization problem in Equation (1) presents challenges related to both uniqueness and computational efficiency. Specifically, the number of vertices $n$ in the thought graph is typically much smaller than the embedding dimension $d$ (i.e., $n\ll d$). For instance, in the GSM8K dataset, thought graphs often contain fewer than 10 vertices, whereas the embedding dimension $d$ can reach up to 768 when using BERT. This large disparity in dimensions leads to a non-unique solution for $W$. Additionally, storing and computing a $d\times d$ matrix is computationally expensive. To address both the issues of uniqueness and computational efficiency, we constrain $W$ to be rank 1, allowing it to be expressed as $W=\alpha\beta^{\top}$. As a result, $\mathcal{A}(W;G)$ simplifies to:
$\mathcal{A}(W;G)=\mathcal{A}(\alpha,\beta;G)=\sum_{i=1}^{n}z_{i}^{\top}\alpha\beta^{\top}x_{i}.$ (3)
Both theoretical analysis and empirical evidence, presented in Appendix A.5, demonstrate that the rank-1 assumption has minimal impact on solving the optimization problem.
Furthermore, with the rank-1 assumption, we can obtain the closed-form solution of Equation (1), as shown in Equation (4), which greatly reduces the computational complexity of solving the optimization problem. The proof is shown in Appendix A.6
$W=\alpha\beta^{\top},\alpha=U[0,:],\beta=V[0,:]$ (4)
$\text{where}\quad U,\Sigma,V=\operatorname{SVD}(X^{\top}Z).$
Discussion. Here, we compare proposed similarity with a widely used attributed-graph similarity to highlight two key properties of our metric: encoding directional relationships and asymmetry.
A common attributed-graph similarity metric computes the graph’s embedding through an iterative formula, and then evaluates the similarity between graphs by calculating the cosine similarity of their embeddings. The iterative formula generally takes the following form:
$X^{(h+1)}=\tilde{A}X^{(h)},\ X^{(0)}=X.$ (5)
Firstly, the embedding-based approach fails to effectively capture the sequential relationships between vertices, which are crucial in multi-step reasoning tasks. In these tasks, the order of operations can significantly alter the reasoning pattern. Therefore, the embedding-based similarity is not well-suited for such tasks. In contrast, our proposed similarity accounts for the transformation of information from previous vertices to current ones, inherently capturing these sequential relationships.
Secondly, our proposed similarity is asymmetric. Specifically, for two thought graphs $G_{1}$ and $G_{2}$, $s(G_{1},G_{2})\neq s(G_{2},G_{1})$. This asymmetry reflects real-world scenarios. For instance, as illustrated in Appendix A.7, mathematical problem A might be a subproblem of problem B. In such a case, referencing B can help resolve A, but referencing A does not necessarily resolve B.
3.3 Example Retrieval
After introducing thought graphs and the proposed similarity metric, we now describe how GraphIC enhances ICL. During the preparation stage, GraphIC generates thought graphs for all candidate examples and estimates their reasoning patterns, denoted as $W_{1},\dots,W_{N}$, where $N$ is the number of candidates. Then, given a query, GraphIC creates a thought graph for the query, denoted as $G^{q}$, computes the applicability of each reasoning pattern to this thought graph, $\mathcal{A}(W_{i},G^{q})$, and then selects the top $k$ as ICEs.
4 Experiments
4.1 Experimental Setup
Datasets and Evaluations. We conduct experiments across four multi-step reasoning tasks: mathematical reasoning (GSM8K Cobbe et al. (2021) and AQUA Ling et al. (2017)), code generation (MBPP Austin et al. (2021)), and logical reasoning (ProofWriter Tafjord et al. (2021)). Model performance is measured by answer accuracy for GSM8K, AQUA, and ProofWriter, and by the pass@1 metric Chen et al. (2021) for MBPP.
Models and Hyper-parameters. We use both an open-source and a closed-source model. For the open-source model, we select Llama-3.1-8B-Instruct (hereafter referred to as Llama-3), one of the most advanced 7B-level models. For the closed-source model, we choose GPT-4o-mini, balancing performance and testing costs. Unless explicitly mentioned otherwise, all evaluations use an 8-shot setting, which is the most common setting in chain-of-thought scenarios. We set the temperature to 1e-5 to minimize randomness, ensuring consistency across generations, and use 3 iterations in Equation (2) with $\lambda$ values from $\{0,0.1,0.2,0.3\}$ (see Appendix 5 for specific details). More experiment details are in Appendix B.1.
4.2 Baselines
We compare GraphIC against six training-free retrieval methods spanning random, similarity-based, diversity-based, and complexity-based approaches, including: (1) Randomrandomly selects $k$ unique ICEs from the candidate set; (2) BM25Robertson et al. (2009)selects the top $k$ examples based on BM25 scoring; (3) BERTDevlin et al. (2019)is a dense retriever using cosine similarity with BERT-base-uncased embeddings; (4) Complex-CoTFu et al. (2022)selects $k$ examples based on complexity, quantified by newline characters; (5) Auto-CoTZhang et al. (2022b)clusters candidates and selects the closests to each cluster center; and (6) Skill-kNNAn et al. (2023)prompts LLM to generate task-relevant skills for query and candidates, followed by dense retrieval.
We also compare with four training-based methods, which encompass both single-example and combination-example retrieval strategies, including: (1) EPRRubin et al. (2022)is trained to retrieve the single most relevant ICE, with top $k$ examples being selected during inference; (2) CEILYe et al. (2023)uses determinantal point processes to select ICEs balancing similarity and diversity; (3) DQ-LoReXiong et al. (2024)uses dual queries and low-rank approximation re-ranking to identify ICEs; and (4) GistScoreGupta et al. (2024)encodes task-specific information into gist tokens for selecting ICEs.
4.3 Main Results
Table 1 shows the performance comparison results. As a training-free method, GraphIC consistently outperforms both training-free and training-based baselines in most settings.
We find that training-based baselines generally outperform training-free baselines, as they learn explicit patterns from in-context examples. For instance, with Llama-3, training-based methods average over 67% performance, while the top training-free baseline (Complex-CoT) reaches only 66.54%. Among the training-based baselines, DQ-LoRe is the most effective, reaching 68.33%. This method leverages LLM-generated outputs during retrieval, demonstrating that incorporating LLM outputs improves performance on reasoning tasks. Among training-free baselines, Complex-CoT and BM25 perform best. Notably, both use asymmetric retrieval, reinforcing that asymmetric retrieval aligns well with real-world reasoning scenarios.
GraphIC integrates the strengths of existing methods—utilizing LLM outputs for reasoning tasks and asymmetric retrieval—while introducing a reasoning-aware representation and similarity metric. Consequently, it not only surpasses all training-free baselines by (2.57% with GPT-4o-mini and 4.29% with Llama-3), but also outperforms all training-based methods (by 1.18% using GPT-4o-mini and 2.5% using Llama-3), highlighting its effectiveness in reasoning tasks.
A closer analysis highlights GraphIC’s substantial advantages in mathematical and logical reasoning tasks compared to code generation, particularly for complex problem instances. For example, on GSM8K, GraphIC outperforms all baselines by 0.65% and 3.57% with the two LLMs. In the more challenging AQUA dataset, the performance improvements become even more pronounced, reaching 3.47% and 7.64%.
LLM Model GSM8K AQUA MBPP ProofWriter Avg.
GPT-4o-mini Random 92.90 (0.31) 71.58 (0.72) 72.76 (0.74) 64.90 (0.93) 75.54 (0.36)
BM25 92.64 70.47 73.4 66.25 75.69
BERT 93.02 66.93 74.2 65.25 74.85
Complex-CoT 92.49 67.32 74.2 64.25 74.57
Auto-CoT 92.72 69.69 73.8 62.25 74.62
Skill-kNN 92.34 71.65 72.0 66.00 75.50
EPR 93.02 72.04 73.8 68.50 76.84
CEIL 92.57 72.44 73.8 69.50 77.08
DQ-LoRe 93.32 69.69 74.6 66.50 76.03
GistScore 93.25 69.69 72.8 67.00 75.69
GraphIC 93.48 73.62 75.2 70.75 78.26
Llama-3.1 -8B-Instruct Random 78.86 (0.87) 53.15 (1.85) 57.72 (1.06) 76.10 (2.45) 66.46 (0.84)
BM25 77.71 46.85 62.0 77.75 66.08
BERT 74.15 50.39 60.8 73.75 64.77
Complex-CoT 79.30 50.00 58.6 78.25 66.54
Auto-CoT 72.78 42.91 58.4 78.00 63.02
Skill-kNN 77.56 50.39 60.8 74.00 65.69
EPR 75.66 53.94 62.0 79.25 67.71
CEIL 75.51 51.97 62.4 81.00 67.72
DQ-LoRe 77.93 54.33 59.8 81.25 68.33
GistScore 74.60 44.49 60.4 79.50 64.75
GraphIC 79.98 57.48 61.6 84.25 70.83
Table 1: Main results on two LLMs and four datasets. For random retrieval, we present the mean and standard deviation derived from five independent experiments.
4.4 Ablation Study
We conduct a series of ablation studies to systematically evaluate the contribution of each component in GraphIC, focusing on three main aspects: filtering shallow semantic information, integrating the graph structure, and the proposed similarity metric.
To this end, we develop several variants of the GraphIC model: (1) Textrelies solely on text embedding, the same as the BERT approach; (2) FRRretrieves examples using text embeddings of FRRs; (3) Graphemploys Equation (5) to generate graph embeddings and use cosine similarity to retrieve examples; (4) Graph+Bemploys Equation (2) to generate graph embeddings and use cosine similarity to retrieve examples; (5) Graph+Sexcludes the backtracking mechanism from the full GraphIC model during computing $Z$; and (6) Graph+B+Srepresents the full GraphIC model, integrating all components.
We conduct experiments using Llama-3 across four datasets, with the results presented in Table 2. First, we observe that employing FRR to compute text embeddings significantly improves performance (FRR versus Text), especially on GSM8K and ProofWriter, with performance increasing from 74.15 to 78.31 and from 73.75 to 82.50, respectively. This improvement is due to FRR’s ability to filter out substantial shallow semantic noise. Second, converting FRR into a graph representation (i.e., a thought graph) further improves performance (Graph versus FRR), as seen on the AQUA dataset, where performance rises from 50.78 to 54.72. This suggests that the graph structure is better suited for capturing the reasoning process than linear text. Third, introducing the reasoning-aware similarity metric also enhances performance (Graph+B+S versus Graph), increasing from 54.72 to 57.48 on AQUA, highlighting the effectiveness of reasoning-aware similarity. Additionally, our experiments show that incorporating the traceback mechanism better simulates the thought process, leading to further performance gains (Graph+B versus Graph, and Graph+B+S versus Graph+S).
In summary, the findings emphasize that each component of GraphIC contributes significantly to improving model performance, with the most substantial gains occurring when all components are combined.
Model GSM8K AQUA MBPP ProofWriter
Text 74.15 50.39 60.8 73.75
FRR 78.31 50.78 60.4 82.50
Graph 78.46 54.72 60.4 83.50
+B 78.92 56.30 61.0 83.75
+S 79.07 49.21 60.4 84.25
+B+S 79.98 57.48 61.6 84.25
Table 2: Ablation Study.
4.5 Analysis
Figure 3: Performance of GraphIC and Top Training-based/Training-free Baselines (DQ-LoRe and Complex-CoT) across 1–8 Shot Settings.
Figure 4: Ground-truth matrix and score matrices of various models. The matrix values have been linearly scaled to the range [0,1], and the diagonal elements have been set to 1.
The performance of GraphIC steadily improves as $k$ increases. We analyze how the performance of GraphIC and top training-based/free baselines (DQ-LoRe and Complex-CoT) changes as $k$ increases. Specifically, we vary the number of examples $k$ in the set {1, 2, 4, 8} and use Llama-3 as LLM.
As shown in Figure 3, we observe that model performance generally improves with $k$. Notably, GraphIC consistently performs better as $k$ increases, whereas other methods experience performance degradation. Additionally, training-based and training-free methods show different trends with $k$ increases. Specifically, training-based methods, such as DQ-LoRe, directly optimize the probability of LLM generating correct answers in 1-shot scenarios. As a result, they tend to achieve superior performance in low-shot settings, especially in 1-shot cases. However, as $k$ increases, the performance gains of these methods tend to slow down or even decline. In contrast, training-free methods, like Complex-CoT and GraphIC, typically underperform compared to training-based methods in low-shot settings. However, as the value of $k$ increases, they exhibit more stable improvement.
Asymmetric retrieval aligns more closely with real-world scenarios. As previously discussed, GraphIC is an asymmetric method that reflects real-world scenarios. To further validate this asymmetry, we randomly select 10 examples from GSM8K and compute their "improve matrix" $S$, which captures the mutual improvement relationships between examples. In this matrix, $S_{ij}$ is defined as the probability of solving the $j$-th example correctly when using the $i$-th example as the ICE. The resulting improve matrix (Figure 4 (a)), clearly illustrates the asymmetry in real-world scenarios: while the $i$-th example may aid in solving the $j$-th example accurately, the reverse is not necessarily true.
Additionally, we examine the alignment between the score matrices of GraphIC and top training-free/based models with the ground-truth improve matrix. Existing ICE retrieval methods select ICEs by assigning scores to (candidate, query) pairs and comparing them. Similar to the improve matrix, we can compute a score matrix for each method, reflecting the mutual improvement relationships between examples estimated by the method. Naturally, the closer a method’s score matrix aligns with the improvement matrix, the better it captures real-world relationships. The score matrices of Complex-CoT, DQ-LoRe, and GraphIC are shown in Figure 4 (b), (c), and (d), respectively. The results clearly demonstrate that GraphIC aligns well with the ground-truth, while the other two models exhibit noticeable deviations.
Additional Analysis. We conducted further experiments to examine other aspects of GraphIC, including the accuracy of the generated pseudo reasoning process (Appendix B.3), comparisons of computational efficiency (Appendix B.4), the impact of the hyperparameter $\lambda$ (Appendix B.5), and the performance on MATH, a more complex mathematical reasoning dataset (Appendxi B.6).
5 Conclusion
We propose GraphIC, which employs a reasoning-aware representation—thought graph and a similarity metric to retrieve ICEs. Our experiments show that GraphIC outperforms both training-free and training-based baselines, highlighting the importance of modeling reasoning structures and providing insights for future research on graph-based solutions for complex problem-solving.
Limitations
Our work has two limitations. First, the GraphIC approach relies on invoking a large language model (LLM) during the retrieval phase, which leads to slower response times compared to embedding-based methods. Second, the thought graph representation and similarity metric used in GraphIC are tailored for multi-step reasoning tasks and may not perform as effectively in other contexts.
Appendix A Supplementary Explanation of GraphIC
A.1 Examples of Thought Graphs
Figure 5 shows examples of thought graphs from the four datasets we considered. For clarity, the label of each graph corresponds to the text associated with its attribute.
Figure 5: Examples of thought graphs for the four datasets we consider. The labels of vertices represent the text for computing embeddings. Dashed-line vertices are included for clarity but are not part of the actual thought graph.
A.2 Formalized Reasoning Representation
The prompt examples below are used to generate formalized reasoning representations for the four datasets being considered. For the test question, since no answer is provided, we will remove the section of the prompt highlighted in red. This will allow the LLM to generate both the answer and the formalized reasoning representation simultaneously, from which we can then extract the formalized reasoning representation.
GSM8K:
Translate the given calculations into code form. Each line of code MUST follow the format specified below:
output_variable = [description of operation](input_variable_1, ..., input_variable_n)
Q: You can buy 4 apples or 1 watermelon for the same price. You bought 36 fruits evenly split between oranges, apples and watermelons, and the price of 1 orange is \$0.50. How much does 1 apple cost if your total bill was \$66?
A: If 36 fruits were evenly split between 3 types of fruits, then I bought 36/3 = <<36/3=12>>12 units of each
fruit
If 1 orange costs \$0.50 then 12 oranges will cost \$0.50 * 12 = \$<<0.5*12=6>>6
If my total bill was \$66 and I spent \$6 on oranges then I spent \$66 - \$6 = \$<<66-6=60>>60 on the other 2 fruit types.
Assuming the price of watermelon is W, and knowing that you can buy 4 apples for the same price and that the price
of one apple is A, then 1W=4A
If we know we bought 12 watermelons and 12 apples for \$60, then we know that \$60 = 12W + 12A
Knowing that 1W=4A, then we can convert the above to \$60 = 12(4A) + 12A
\$60 = 48A + 12A
\$60 = <<60=60>>60A
Then we know the price of one apple (A) is \$60/60= \$<<60/60=1>>1
#### 1
Code:
total_fruits = 36
types_of_fruits = 3
price_per_orange = 0.50
total_oranges = 12
total_bill = 66
equivalent_apples_for_watermelon = 4
total_apples_and_watermelons = 12
fruits_per_type = [divide](total_fruits, types_of_fruits)
cost_of_oranges = [multiply](total_oranges, price_per_orange)
remaining_budget = [minus](total_bill, cost_of_oranges)
price_per_apple = [construct and solve an equation](total_apples_and_watermelons, equivalent_apples_for_watermelon, remaining_budget)
Q: {{question}}
A: {{answer}}
Code:
AQUA:
Translate the given calculations into code form. Each line of code MUST follow the format specified below:
output_variable = [description of operation](input_variable_1, ..., input_variable_n)
Q: In a group of 6 boys and 4 girls, four children are to be selected. In how many different ways can they be selected such that at least one boy should be there?
Options: A)209, B)210, C)211, D)213, E)215
A: To determine the number of ways to select 4 children from a group of 6 boys and 4 girls such that at least one boy is included, we will use the method of complement counting.
First, lets calculate the total number of ways to select 4 children from 10 children (6 boys + 4 girls):
[…]
Next, we calculate the number of ways to select 4 children with no boys, i.e., all girls. Since there are only 4 girls, and we need to select all 4 of them:
[…]
Now, subtract the number of ways to select all girls from the total number of ways to select 4 children to find the number of ways that include at least one boy:
[…]
Thus, the number of ways to select 4 children with at least one boy is:
[…]
#### A
Code:
total_children = 10
children_to_select = 4
boys = 6
girls = 4
total_ways_to_select = [combination](total_children, children_to_select)
all_girls_selection = [combination](girls, children_to_select)
ways_with_at_least_one_boy = [subtract](total_ways_to_select, all_girls_selection)
Q: {{question}}
A: {{answer}}
Code:
MBPP:
Text: Write a function to generate a square matrix filled with elements from 1 to n raised to the power of 2 in spiral order.
Test Cases:
assert generate_matrix(3)==[[1, 2, 3], [8, 9, 4], [7, 6, 5]]
assert generate_matrix(2)==[[1,2],[4,3]]
assert generate_matrix(7)==[[1, 2, 3, 4, 5, 6, 7], [24, 25, 26, 27, 28, 29, 8], [23, 40, 41, 42, 43, 30, 9], [22, 39, 48, 49, 44, 31, 10], [21, 38, 47, 46, 45, 32, 11], [20, 37, 36, 35, 34, 33, 12], [19, 18, 17, 16, 15, 14, 13]]
Code:
def generate_matrix(n):
if n<=0:
return []
matrix=[row[:] for row in [[0]*n]*n]
row_st=0
row_ed=n-1
col_st=0
col_ed=n-1
current=1
while (True):
if current>n*n:
break
for c in range (col_st, col_ed+1):
matrix[row_st][c]=current
current+=1
row_st+=1
for r in range (row_st, row_ed+1):
matrix[r][col_ed]=current
current+=1
col_ed-=1
for c in range (col_ed, col_st-1, -1):
matrix[row_ed][c]=current
current+=1
row_ed-=1
for r in range (row_ed, row_st-1, -1):
matrix[r][col_st]=current
current+=1
col_st+=1
return matrix
Text: {{question}}
Test Cases:
{{test_list}}
Code:
ProofWriter:
Formalize the following proof process based on the given information. Where Each line of formalized proof MUST follow the format specified below.
output_triple = [rule used](input_triple_1, ... , input_triple_n)
Q: Triples:
1. Anne is not big.
2. Anne is cold.
3. Anne is red.
4. Dave is green.
5. Dave is rough.
6. Erin is green.
7. Erin is kind.
8. Erin is rough.
9. Fiona is green.
10. Fiona is not nice.
Rules:
1. If Erin is cold then Erin is rough.
2. If something is rough then it is nice.
3. All green, big things are kind.
4. If Dave is kind then Dave is cold.
5. If something is green and not rough then it is big.
6. All nice, rough things are big.
7. If Dave is cold and Dave is nice then Dave is red.
Based on the above information, is the following statement true or false? Dave is red.
A) True B) False
Proof: To determine if the statement "Dave is red" is true or false, lets evaluate the given information step by step:
From the triples:
1. Dave is green.
2. Dave is rough.
From the rules:
1. If something is rough then it is nice. (Dave is rough, so Dave is nice.)
2. If Dave is cold and Dave is nice then Dave is red. (However, theres no information saying Dave is cold, so we cant use this rule directly.)
3. All nice, rough things are big. (Dave is nice and rough, so Dave is big.)
Now, we need to verify whether Dave being cold can be derived:
1. If Dave is kind then Dave is cold. (To use this, we need to verify if Dave is kind.)
- All green, big things are kind. (Dave is green and big, so Dave is kind.)
- Hence, Dave is kind, which according to rule 4, makes Dave cold.
Since now we know:
- Dave is cold.
- Dave is nice.
We can use rule 7:
- If Dave is cold and Dave is nice then Dave is red.
Since both conditions are satisfied, Dave is red.
The statement "Dave is red" is **True (A)**.
Formalized Proof:
"Dave is nice." = [rule2]("Dave is rough.")
"Dave is big." = [rule6]("Dave is nice.", "Dave is rough.")
"Dave is kind." = [rule3]("Dave is green.", "Dave is big.")
"Dave is cold." = [rule4]("Dave is kind.")
"Dave is red." = [rule7]("Dave is cold.", "Dave is nice.")
Proof: {{answer}}
Formalized Proof:
The pseudo-code of a parser that transforms formalized reasoning representations into a thought graph is provided in Algorithm 1. Inputs, Output, and OperationName are extracted following the pattern outlined below.
Output = [OperationName](input_1, ..., input_n)
Algorithm 1 Parsing Formalized Reasoning Representation
0:  formalized reasoning representation $FRR$
0:  Corresponding graph $G(V,E)$
[…]
3:  $line\leftarrow$ first line of $FRR$
4:  while $line\neq$ NULL do
5:     Extract Inputs, Output, and OperationName from $line$
6:     for each $input$ in Inputs do
7:        if $input\notin\textit{NodeSet}$ then
8:           Add $input$ to $V$
9:        end if
10:     end for
11:     if Output $\notin V$ then
12:        Add Output to $V$, labeled as OperationName
13:     end if
14:     for each $input$ in Inputs do
15:        Add directed edge from $input$ to Output to $E$
16:     end for
17:     $line\leftarrow$ next line of $FRR$
18:  end while
19:  $G=G(V,E)$
A.3 Robustness to the Correctness of Pseudo Reasoning Process
As mentioned in Section 3.1, ensuring the correctness of the generated pseudo-reasoning process is not required. However, we have the following reasons for ensuring that GraphIC maintains high performance even when the generated process is incorrect.
  1. 1.
    Firstly, even if the generated answers are not correct, it does not necessarily mean that the reasoning steps are entirely incorrect; rather, many are "almost correct," involving only minor errors Wei et al. (2022). This is also helpful for retrieving examples with similar reasoning processes.
  2. 2.
    To further illustrate this fact, we selected a subset from each dataset, containing all queries associated with incorrect pseudo answers. We evaluated GraphIC on these four subsets and compared its performance with top training-based (DQ-LoRe) and training-free (Complex-CoT) baselines. The results, shown in Table 3, reveal that even when GraphIC uses incorrect thought graphs to retrieve in-context examples, it still achieves a significant performance advantage. Note that the performance in Table 3 are substantially lower than those in Table 1. This is because the queries that lead to incorrect thought graphs are typically the most difficult ones.
  3. 3.
    Similar approaches have been applied in other ICE retrieval models and show good performance, such as Skill-kNN, which utilizes an LLM to generate the skills required to solve a problem, and DQ-LoRe, which uses an LLM to generate chain-of-thought reasoning answers.
Model GSM8K AQUA MBPP ProofWriter
Complex-CoT 38.58 28.68 4.06 54.02
DQ-LoRe 40.83 33.33 16.75 65.51
GraphIC 43.08 33.33 15.23 70.11
Table 3: Performance of GraphIC and top training-free/based baseline on the subset where GraphIC uses incorrect answers for creating thought graphs.
A.4 The Matrix $B$
As shown in Figure 6, matrix $A$ represents the adjacency matrix of a thought graph, corresponding to the black edges in the graph. In this graph, vertices 1 and 2 have an in-degree of 0, indicating the starting points of reasoning, while vertices 3 and 4 have non-zero in-degrees, representing intermediate steps or results of reasoning. Matrix $B$ indicates the edges from vertices with non-zero in-degrees to those with in-degree 0, corresponding to the edges from vertices 3 and 4 to vertices 1 and 2 in the graph (marked in green). Since vertices 3 and 4 are intermediate steps or results of reasoning, and vertices 1 and 2 represent the starting points, these edges represent the retracing process.
Figure 6: The definition of matrix $B$ and its corresponding edges in a graph.
A.5 The Loss of Setting the Rank of Matrix $W$ to 1
First, we provide the optimal value and the optimal solution of the optimization problem defined by Equation (1), under the assumption that no constraints are imposed on $W$.
Theorem 1
Consider the following optimization problem:
$\mathop{\max}\limits_{\begin{subarray}{c}W\in\mathbb{R}^{d\times d},\\ \|W\|_{F}=1\end{subarray}}\sum_{i=1}^{n}z_{i}^{\top}Wx_{i}.$ (6)
The optimal value of this problem is $\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}$, and the optimal solution is $W^{*}=VYU^{\top}$, where $Y=\text{diag}(y_{11},y_{22},\dots,y_{dd})$ with $y_{ii}=\frac{\sigma_{i}}{\left(\sum_{j=1}^{d}\sigma_{j}^{2}\right)^{1/2}}$.
First, the objective function $\sum_{i=1}^{n}z_{i}^{\top}Wx_{i}$ can be written as $\operatorname{tr}(ZW^{\top}X^{\top})$
Then, we rewrite the objective function using the cyclic property of the trace:
$\text{tr}(ZW^{\top}X^{\top})=\text{tr}(W^{\top}X^{\top}Z).$
Next, perform the Singular Value Decomposition (SVD) of $X^{\top}Z$:
$X^{\top}Z=U\Sigma V^{\top},$
where $U$ and $V$ are orthogonal matrices, and $\Sigma=\text{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{d})$ is the diagonal matrix of singular values.
Substituting this decomposition into the trace expression, we obtain:
$\text{tr}(ZW^{\top}X^{\top})=\text{tr}(W^{\top}U\Sigma V^{\top}).$
Using the cyclic property of the trace again, we get:
$\text{tr}(ZW^{\top}X^{\top})=\text{tr}\left((V^{\top}W^{\top}U)\Sigma\right).$
Define $Y=V^{\top}W^{\top}U$. Then the objective becomes:
$\text{tr}(ZW^{\top}X^{\top})=\text{tr}(Y\Sigma).$
Since $U$ and $V$ are orthogonal, they preserve the Frobenius norm, i.e., $||Y||_{F}=||W||_{F}=1$. Thus, we are now tasked with maximizing $\text{tr}(Y\Sigma)$ subject to $||Y||_{F}=1$.
We know that the trace $\text{tr}(Y\Sigma)$ is maximized when $Y$ is diagonal, i.e., when $Y=\text{diag}(y_{11},y_{22},\dots,y_{dd})$. This follows because the off-diagonal elements of $Y$ do not contribute to $\text{tr}(Y\Sigma)$, but they affect the Frobenius norm of $Y$. By setting these off-diagonal elements to zero and redistributing the weight to the diagonal elements, we achieve a higher value of $\text{tr}(Y\Sigma)$.
Thus, the optimization problem reduces to:
$\text{tr}(ZW^{\top}X^{\top})=\sum_{i=1}^{d}\sigma_{i}y_{ii},\ \text{s.t.}\ \sum_{i=1}^{d}y_{ii}^{2}=1.$
To find the optimal $y_{ii}$, we apply Cauchy-Schwarz inequality:
$\sum_{i=1}^{d}\sigma_{i}y_{ii}\leq(\sum_{i=1}^{d}\sigma_{i}^{2})^{1/2}(\sum_{i=1}^{d}y_{ii}^{2})^{1/2}.$
Since $\sum_{i=1}^{d}y_{ii}^{2}=1$, this simplifies to:
$\sum_{i=1}^{d}\sigma_{i}y_{ii}\leq(\sum_{i=1}^{d}\sigma_{i}^{2})^{1/2}.$
The maximum value of $\sum_{i=1}^{d}\sigma_{i}y_{ii}$ is achieved when $y_{ii}=\frac{\sigma_{i}}{\left(\sum_{j=1}^{d}\sigma_{j}^{2}\right)^{1/2}}$. Thus, the optimal value of the objective is $\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}$, and the corresponding optimal solution is $W^{*}=VYU^{\top}$, where $Y=\text{diag}\left(\frac{\sigma_{1}}{\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}},\dots,\frac{\sigma_{d}}{\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}}\right)$.
Based on the proof above, we know that the optimal value of the optimization problem defined by Equation (1) is $\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}$. When a rank-1 constraint is added to $W$, the optimal value of the problem becomes $\sigma_{1}$. Therefore, we are interested in quantifying the loss introduced by the "rank-1 assumption," which can be assessed by the ratio of their optimal values, $r=\sigma_{1}/\left(\sum_{i=1}^{d}\sigma_{i}^{2}\right)^{1/2}$. We computed the value of $r$ on four datasets, and the results are shown in Table 4. The table demonstrates that the loss caused by the "rank-1 assumption" is less than 0.1%, implying that it does not result in a significant loss of precision.
Dataset GSM8K AQUA MBPP ProofWriter
$r$ 0.99978 0.99991 0.99963 0.99921
Table 4: The $r$ value on four datasets.
A.6 Closed Form Solution of Equation (1) with Rank-1 Assumption
Under the rank-1 assumption, the optimization problem can be expressed as
$\mathop{\max}\limits_{\begin{subarray}{c}\alpha,\beta\in\mathbb{R}^{d},\\ \|\alpha\|_{2}=\|\beta\|_{2}=1\end{subarray}}\sum_{i=1}^{n}z_{i}^{\top}\alpha\beta^{\top}x_{i}.$
This can be rewritten using the trace operator as
$\operatorname{tr}(Z\beta\alpha^{\top}X^{\top})=\operatorname{tr}(\alpha^{\top}X^{\top}Z\beta)=\alpha^{\top}X^{\top}Z\beta.$
Thus, the problem becomes a bilinear maximization:
$\mathop{\max}\limits_{\begin{subarray}{c}\alpha,\beta\in\mathbb{R}^{d},\\ \|\alpha\|_{2}=\|\beta\|_{2}=1\end{subarray}}\alpha^{\top}X^{\top}Z\beta.$
The closed-form solution to this problem, as derived by Leon (1994), is:
$\alpha=U[0,:],\quad\beta=V[0,:],$
where $U,\Sigma,V$ are the singular value decomposition (SVD) of $X^{\top}Z$.
A.7 The Asymmetry
We provide an example to illustrate the asymmetry. As shown in the example below, solving Question B includes solving Question A, which involves further calculations to determine how long it will take to reach the minimum age required by the company for employment. Therefore, in this case, referencing B can help resolve A, but referencing A does not necessarily resolve B.
Question A:
In 3 years, Jayden will be half of Ernestos age. If Ernesto is 11 years old, how many years old is Jayden now?
Answer A:
Ernesto = 11 + 3 = 14
Jayden = 14/2 = 7 in 3 years
Now = 7 - 3 = 4
Jayden is 4 years old.
Question B:
The minimum age required to be employed at a company is 25 years. Dara aspires to work for the company and will be half the age of Jane in six years. If Jane is currently working for the company and is 28 years old, how long is it before Dara reaches the minimum age required by the company to be employed?
Answer B:
In six years, Jane will be 28+6 = 34 years old.
Dara will be half the age of Jane in six years, meaning she will be 34/2 = 17 years old in six years.
Currently, Dara is 17-6 = 11 years old.
Dara has to wait for 25-11 = 14 more years to reach the companys minimum age of employment.
Appendix B Supplementary Materials of Experiments
B.1 Further Experiment Details
Dataset Preprocessing.
For both GSM8K and MBPP, we utilize the original datasets without further preprocessing. For AQUA and ProofWriter, we refine the original dataset to improve the experimental setup, as described below.
Given the substantial size of the AQUA dataset, which incurs significant retrieval overhead during testing, we followed the methodology outlined in DQ-LoRe Xiong et al. (2024), using a 1,000-sample subset for efficient evaluation.
For the ProofWriter dataset, we refined the subset selected by Logic-LM Pan et al. (2023), excluding instances labeled as “Unknown," as these samples lacked explicit reasoning chains. Furthermore, because the original training set did not provide reasoning in natural language, we leveraged the GPT-4o-mini model to generate reasoning sequences for the training set, discarding any generated outputs deemed incorrect. We evaluate the correctness of the reasoning process by the correctness of the final result, which is a commonly used approach Lightman et al. (2024); Xiong et al. (2024); Khattab et al. (2022). This process resulted in a refined training set of 1,358 examples with their Chains of Thought and 400 test samples from the original ProofWriter dataset.
Thought Graph Construction and Embedding Model.
For GSM8K, AQUA, and ProofWriter, we prompt the LLM to create a FRR for thought graph construction, using vertex features from a BERT model. For MBPP, we use the staticfg [^1] https://github.com/coetaur0/staticfg module to parse Python code and generate the control flow graph, embedding each vertex’s features with CodeBERT Feng et al. (2020).
B.2 Values of $\lambda$
We select hyper parameter $\lambda$ values from $\{0,0.1,0.2,0.3\}$, and report the $\lambda$ values chosen on various datasets and LLMs in Table 5.
Engine GSM8K AQUA MBPP ProofWriter
GPT-4o-mini 0.2 0.2 0.1 0.1
Llama-3 0.3 0.2 0.2 0.0
GPT-3.5-Turbo 0.3 / / /
Table 5: $\lambda$ values chosen on various datasets and LLMs.
B.3 Correctness of LLM Generated Answers for Creating Thought Graphs
To analyze the consistency between LLM-generated answers and real solutions, we tested the accuracy of these answers used to generate the thought graphs. The results are shown in the Table 6. From Table 6, it can be seen that these answers used to generate the thought graph already have relatively high accuracy, which ensures their consistency with the real solution. Furthermore, the table demonstrates that using the thought graph to retrieve examples can further improve accuracy, especially in mathematical reasoning and logical reasoning tasks. We use Llama-3.1-8B-Instruct for testing.
Accuracy GSM8K AQUA MBPP ProofWriter
Pseudo Answer 76.42 49.21 60.6 78.25
Final Answers 79.98 57.48 61.6 84.25
Improvement +3.56 +8.27 +1.00 +6.00
Table 6: Correctness of LLM generated answers for creating thought graphs and final answers.
B.4 Comparison of Computation Time with Other Similar Baselines.
Our method belongs to the category of methods that use the generated output of LLMs to select in-context examples, which also includes methods such as Skill-kNN and DQ-LoRe. These approaches involve using the LLM during the retrieval phase, resulting in longer retrieval times compared to other baselines. However, by leveraging the power of LLMs, they are suitable for complex tasks. The computation times for the three models are presented in Table 7. Specifically, the retrieval time for the GraphIC model is similar to that of DQ-LoRe, and slightly higher than Skill-kNN. Despite this, GraphIC significantly outperforms Skill-kNN in terms of performance. Moreover, compared to DQ-LoRe, which has the same retrieval time, GraphIC not only delivers superior performance but also greatly reduces both the prepare time and training time required by DQ-LoRe.
Here, "prepare time" refers to the time spent generating the necessary outputs for retrieval, such as generating the required skills for all candidate examples in Skill-kNN. For our evaluation, we used the GSM8K dataset with the LLM configured as Llama-3.1-8B-Instruct.
Time Skill-kNN DQ-LoRe GraphIC
prepare time 0.7h 20h 1.5h
traning time - 16h -
retrieve time 0.3s 0.4s 0.4s
Table 7: Prepare time, training time, and retrieve time of GraphIC and other similar baselines.
B.5 The Effects of $\lambda$
We analyzed the effect of $\lambda$ values ranging from [0.0, 0.9] on the results across the four datasets utilized in our study. The corresponding results are presented in Figure 7, confirming the robustness of our method to the choice of lambda.
Figure 7: The performance of GraphIC on different $\lambda$ across four datasets.
B.6 The Performance on MATH Dataset
We conducted a comparison of GraphIC with the top-performing training-based and training-free baselines (DQ-LoRe and Complex-CoT) on MATH Hendrycks et al. (2021) dataset. The results are presented in Table 8: the GraphIC model has consistently achieved optimal performance. We randomly selected 500 samples from the training and testing data categorized as "level 5" difficulty in MATH, which were used as the candidate set and test set.
Model Complex-CoT DQ-LoRe GraphIC
Llama-3 18.8 20.4 21.8
GPT-4o-mini 58.8 59.6 61.2
Table 8: The Performance on MATH Dataset
B.7 The Performance on GSM8K Using GPT-3.5-Turbo
Random BM25 BERT Complex-CoT Auto-CoT Skill-kNN EPR CEIL DQ-LoRe GistScore GraphIC
80.76(0.55) 82.10 80.89 81.65 82.03 81.50 81.65 81.72 82.10 81.72 82.79
Table 9: Results obtained using GPT-3.5-Turbo as the LLM on the GSM8K dataset.
We observe that the GPT-4o-mini model’s performance on the GSM8K dataset is relatively invariant to the selection of ICEs. So, we further perform an additional experiment on the GSM8K dataset using the GPT-3.5-Turbo model. As presented in Table 9, GraphIC achieves superior performance across all metrics.